### v2

@@ IV-B2 Effectiveness of derived inequalities
To illustrate the effectiveness of inequalities (26)–(31) in strengthening the FP problem, let us consider a special case in which $\ell_{k}=1$ for all $k\in\mathcal{K}$ in [36]. In this case, the nonlinear terms $\sum_{s\in\mathcal{F}^{k}\backslash\{\ell_{k}\}}\lambda_{s}^{k}{(x^{k,s+1}_{v}-x^{k,s}_{v})}^{+}$ and $\sum_{s\in\mathcal{F}^{k}\backslash\{\ell_{k}\}}\lambda_{s}^{k}{(x^{k,s}_{v}-x^{k,s+1}_{v})}^{+}$ in (23) and (25) can be removed, and thus we do not need to introduce variables $\{z_{v}^{k,s}\}$ and $\{w_{v}^{k,s}\}$. Moreover, for this case, inequalities (3), (23), and (25) can be combined into
@@
$x_{B}^{k,1}+x_{B}^{k,2}\leq\min\left\{\mu_{B},C_{AB},C_{BD}\right\}y_{B}=y_{B},$ (33)
$x_{C}^{k,1}+x_{C}^{k,2}\leq\min\left\{\mu_{C},C_{AC},C_{CD}\right\}y_{C}=y_{C},$ (34)
will yield a solution with an objective value being $3$ (in which both cloud nodes are activated, as to process the functions of the two services). This example clearly shows that adding constraints (32) (or (26)–(31) for the general case) can effectively strengthen the FP problem in terms of providing a better relaxation bound of problem (NS).
The strongness of relaxation (FP-II) makes it much more suitable to be embedded into the proposed decomposition algorithm in Section III than relaxation (FP-I). In particular, with this stronger relaxation (FP-II), the proposed decomposition algorithm converges much faster, thereby enjoying a much better overall performance; see Section V-A further ahead.
@@ V Numerical Results
In this section, we present numerical results to demonstrate the effectiveness and efficiency of the proposed valid inequalities and decomposition algorithm. More specifically, we first perform numerical experiments to illustrate the effectiveness of the proposed valid inequalities for the proposed decomposition algorithm in Section V-A. Then, in Section V-B, we present numerical results to demonstrate the effectiveness and efficiency of the proposed decomposition algorithm over the SOTA approaches in [7, 16, 17].
@@ V-A Effectiveness of Proposed Valid Inequalities
In this subsection, we illustrate the effectiveness of our proposed connectivity based inequalities (17) and (19) and link-capacity based inequalities (26)–(31). To do this, we first compare the optimal values of relaxations (FP), (FP-I), and (FP-II). We compare the relative optimality gap improvement, defined by
@@
where $\text{x}\in\{\text{I},\text{II}\}$ and $\nu(\cdot)$ denotes the optimal objective value of the corresponding problem. The gap improvement in (35) quantifies the improved tightness of relaxations (FP-I) and (FP-II) over that of relaxation (FP), and thus the effectiveness of proposed connectivity based inequalities (17) and (19), and link-capacity based inequalities (26)–(31) in reducing the gap between the FP problem and the NS problem. The larger the gap improvement, the stronger the relaxations (FP-I) and (FP-II) (as compared with relaxation (FP)), and the more effective the proposed inequalities. Note that the gap improvement is a widely used performance measure in the integer programming community [37, 38] to show the tightness of a relaxation problem over another one.
@@
Figure 7: The average number of iterations needed by DECOs with relaxations (FP), (FP-I), and (FP-II), respectively.
@@
Figure 8: The average CPU time taken by DECOs with relaxations (FP), (FP-I), and (FP-II), respectively.
@@
Fig. 6 plots the average gap improvement versus different numbers of services. As can be observed in Fig. 6, in all cases, the gap improvement is larger than $0.1$, which shows that (FP-I) and (FP-II) are indeed stronger than (FP). Compared with that of (FP-I), the gap improvement of (FP-II) is much larger. This indicates that the link-capacity based inequalities (26)–(31) can indeed significantly strengthening the FP problem only with the connectivity based inequalities (17) and (19). In addition, with the increasing of the number of services, the gap improvement of (FP-II) becomes larger. This can be explained as follows. As the number of services increases, the left-hand sides of inequalities (30) and (31) become larger, thereby making inequalities (30) and (31) tighter. Consequently, the gap improvement achieved by the link-capacity based inequalities is also likely to be larger.
@@
Next, we illustrate the efficiency of the proposed connectivity based inequalities (17) and (19) and link-capacity based inequalities (26)–(31) for accelerating the convergence of the decomposition algorithm (called DECO). Figs. 7 and 8 plot the average numbers of iterations and CPU times needed by the proposed DECOs with relaxations (FP), (FP-I), and (FP-II), respectively, to terminate. As observed, both of the derived families of valid inequalities can improve the performance of the proposed DECO. In particular, with the two proposed inequalities, the average number of iterations needed for the convergence of DECO is much smaller (less than $2$); and thus the CPU time taken by DECO is also much smaller.
@@
From the above computational results, we can conclude that the proposed DECO with relaxation (FP-II) (i.e., the FP problem with the two families of valid inequalities) significantly outperforms that with relaxation (FP) (i.e., the FP problem without the two families of valid inequalities). Due to this, we shall only test and compare the proposed DECO with relaxation (FP-II) with the other existing algorithms in the following.
V-B Comparison of DECO with SOTA Algorithms
@@
Figure 9: The number of feasible problem instances solved by DECOs with $\text{IterMax}=5,\,+\infty$, LPR, PSUM-R, and MBLP-S.
@@
Figure 10: The average objective value returned by DECOs with $\text{IterMax}=5,\,+\infty$, LPR, PSUM-R, and MBLP-S.
@@
Figure 11: The average CPU time returned by DECOs with $\text{IterMax}=5,\,+\infty$, LPR, PSUM-R, and MBLP-S.
@@
In this subsection, we compare the performance of the proposed DECO with SOTA algorithms such as the exact approach using standard MBLP solvers (called MBLP-S) in [7], the LP rounding (LPR) algorithm in [16], and the penalty successive upper bound minimization with rounding (PSUM-R) algorithm in [17].
@@
Figs. 9 and 10 plot the number of feasible problem instances and the average objective value returned by the proposed DECOs with $\text{IterMax}=5,\,+\infty$, LPR, PSUM-R, and MBLP-S, respectively. When setting $\text{IterMax}=5$, the proposed DECO cannot be guaranteed to find an optimal solution of problem (NS), and the red-diamond curve in Fig. 9 is obtained as follows: for a problem instance, if the proposed DECO can find an optimal solution of problem (NS) within 5 iterations, it is counted as a feasible instance; otherwise, it is infeasible. From Figs. 9 and 10, we can observe the effectiveness of the proposed DECOs over LPR and PSUM-R. In particular, compared with LPR and PSUM-R, the proposed DECOs can find a feasible solution for much more problem instances and return a much better solution with a much smaller objective value. The latter is reasonable, as stated in Theorem 3, DECOs must return an optimal solution as long as they find a feasible solution. Another simulation finding is that with the increasing of IterMax, DECO can return a feasible solution for more problem instances. Nevertheless, when $\text{IterMax}=5$, DECO can return a feasible solution for almost all instances (as DECO with $\text{IterMax}=+\infty$ is able to find all truly feasible problem instances and the difference between the numbers of feasible problem instances solved by DECOs with $\text{IterMax}=5$ and $\text{IterMax}=+\infty$ is small in Fig. 9). This shows that in most cases, DECO can return an optimal solution within a few number of iterations, thanks to the two derived families of valid inequalities.
@@
The comparison of the solution efficiency of the proposed DECOs with $\text{IterMax}=5,\,+\infty$, LPR, PSUM-R, and MBLP-S is plotted in Fig. 11. We can observe from Fig. 11 that though both MBLP-S and the proposed DECO with $\text{IterMax}=+\infty$ can return an optimal solution (if the problem is feasible), DECO with $\text{IterMax}=+\infty$ is much more computationally efficient than MBLP-S. In particular, the average CPU times taken by DECO with $\text{IterMax}=+\infty$ are less than $0.7$ seconds, while the CPU times taken by MBLP-S in the large-scale cases (i.e., $|\mathcal{K}|\geq 13$) are generally larger than $5$ seconds. The solution efficiency of the proposed DECO with $\text{IterMax}=5$, LPR, and PSUM-R is comparable. In general, LPR performs the best in terms of the CPU time, followed by DECO with $\text{IterMax}=5$ and PSUM-R.
@@
In summary, our simulation results show both of the effectiveness and efficiency of the proposed DECO. More specifically, compared with the LPR and PSUM-R algorithms in [16] and [17], the proposed DECO is able to find a much better solution (due to the information feedback between the two subproblems); compared with the exact approach in [7], the proposed DECO is significantly more computationally efficient, which is because of its decomposition nature.
@@ VI Conclusions
In this paper, we have proposed an efficient decomposition algorithm for solving large-scale NS problems. The proposed algorithm decomposes the large-scale NS problem into two relatively easy FP and TR subproblems and solves them in an iterative fashion enabling a useful information feedback between them (to deal with the coupled constraints) until an optimal solution of the problem is found. To further improve the efficiency of the proposed decomposition algorithm, we have also developed two families of valid inequalities to speed up its convergence by judiciously taking the special structure of the considered problem into account. Two key features of the proposed algorithm, which make it particularly suitable to solve the large-scale NS problems, are as follows. First, it is guaranteed to find an optimal solution of the NS problem if the number of iterations between the two subproblems is allowed to be sufficiently large and it can return a much better solution than existing two-stage heuristic algorithms even though the number of iterations between the subproblems is small (e.g., 5). Second, the FP and TR subproblems in the proposed decomposition algorithm are a small MBLP problem and a polynomial time solvable LP problem, respectively, which are much easier to solve than the original problem. Simulation results show the effectiveness and efficiency of the proposed decomposition algorithm over the existing SODA algorithms in [7, 16, 17]. In particular, when compared with algorithms based on LP relaxations [16, 17], the proposed decomposition algorithm is able to find a much better solution; when compared with the algorithm needed to call MBLP solvers [7], the proposed decomposition algorithm is much more computationally efficient.
@@ Appendix A Proof of Theorem 5
$x_{v_{0}}^{k,s_{0}}+x_{v}^{k,s_{0}+1}\leq 1,~{}\forall~{}v_{0}\in\mathcal{V}\backslash\mathcal{V}(S^{k}),~{}\forall~{}v\in\mathcal{V}(v_{0}).$ (40)
@@
Next, we show that (17) holds for all $k\in\mathcal{K}$, $v\in\mathcal{V}(D^{k})$, and $s\in\mathcal{F}^{k}$ at $\bm{x}\in\mathcal{X}^{1}$ by an induction of $s$. By (15), $x_{v}^{k,s}=0$ holds for $s=\ell_{k}$. Suppose that $x_{v}^{k,s}=0$ holds for $s=s_{0},s_{0}+1,\ldots,\ell_{k}$, where $1<s_{0}\leq\ell_{k}$. Then we have $x_{v}^{k,s_{0}}=0$ for all $v\in\mathcal{V}(D^{k})$. This, together with (1), implies
@@
$x_{v_{0}}^{k,s_{0}-1}+x_{v}^{k,s_{0}}\leq 1,~{}\forall~{}v_{0}\in\mathcal{V}(D^{k}),~{}\forall~{}v\in\mathcal{V}(v_{0}).$ (43)

### v3

@@ IV-B2 Effectiveness of derived inequalities
To illustrate the effectiveness of inequalities (24) and (26) (or their linearization (28), (29), (32), (33), (35), and (36)) in strengthening the FP problem, let us consider a special case in which $\ell_{k}=1$ for all $k\in\mathcal{K}$ in [40]. In this case, the nonlinear terms $\sum_{s\in\mathcal{F}^{k}\backslash\{\ell_{k}\}}\lambda_{s}^{k}{(x^{k,s+1}_{v}-x^{k,s}_{v})}^{+}$ and $\sum_{s\in\mathcal{F}^{k}\backslash\{\ell_{k}\}}\lambda_{s}^{k}{(x^{k,s}_{v}-x^{k,s+1}_{v})}^{+}$ in (24) and (26) can be removed, and thus we do not need to introduce variables $\{z_{v}^{k,s}\}$ and $\{w_{v}^{k,s}\}$. Moreover, for this case, inequalities (3), (24), and (26) can be combined into
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{x_{B}^{1}+x_{B}^{2}}}\leq{\min}\left\{\mu_{B},C_{AB},C_{BD}\right\}y_{B}=y_{B},$ (38)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{x_{C}^{1}+x_{C}^{2}}}\leq{\min}\left\{\mu_{C},C_{AC},C_{CD}\right\}y_{C}=y_{C},$ (39)
will yield a solution with an objective value being $3$ (in which both cloud nodes are activated, as to process the functions of the two services). This example clearly shows that adding constraints (37) (or (28), (29), (32), (33), (35), and (36) for the general case) can effectively strengthen the FP problem in terms of providing a better relaxation bound of problem (NS).
The strongness of relaxation (FP-II) makes it much more suitable to be embedded into the proposed Algorithm 1 than relaxation (FP-I). In particular, with this stronger relaxation (FP-II), the proposed Algorithm 1 converges much faster, thereby enjoying a much better overall performance; see Section V-A further ahead.
@@ V Numerical Results
In this section, we present numerical results to demonstrate the effectiveness and efficiency of the proposed valid inequalities and decomposition algorithm. More specifically, we first perform numerical experiments to illustrate the effectiveness of the proposed valid inequalities for the proposed decomposition algorithm in Section V-A. Then, in Section V-B, we present numerical results to demonstrate the effectiveness and efficiency of the proposed CBD algorithm over the SOTA approaches in [7, 16, 17, 18].
@@ V-A Effectiveness of Proposed Valid Inequalities
In this subsection, we illustrate the effectiveness of our proposed connectivity based inequalities (18) and (20) and link-capacity based inequalities (28), (29), (32), (33), (35), and (36). To do this, we first compare the optimal values of relaxations (FP), (FP-I), and (FP-II). We compare the relative optimality gap improvement, defined by
@@
where $\text{x}\in\{\text{I},\text{II}\}$ and $\nu(\cdot)$ denotes the optimal objective value of the corresponding problem. The gap improvement in (40) quantifies the improved tightness of relaxations (FP-I) and (FP-II) over that of relaxation (FP), and thus the effectiveness of proposed connectivity based inequalities (18) and (20), and link-capacity based inequalities (28), (29), (32), (33), (35), and (36) in reducing the gap between the FP problem and the NS problem. The larger the gap improvement, the stronger the relaxations (FP-I) and (FP-II) (as compared with relaxation (FP)), and the more effective the proposed inequalities. Note that the gap improvement is a widely used performance measure in the integer programming community [41, 42] to show the tightness of a relaxation problem over another one.
@@
Figure 6: The average number of iterations needed by CBDs with relaxations (FP), (FP-I), and (FP-II), respectively.
@@
Figure 7: The average CPU time taken by CBDs with relaxations (FP), (FP-I), and (FP-II), respectively.
@@
Fig. 5 plots the average gap improvement versus different numbers of services. As can be observed in Fig. 5, in all cases, the gap improvement is larger than $0.1$, which shows that (FP-I) and (FP-II) are indeed stronger than (FP). Compared with that of (FP-I), the gap improvement of (FP-II) is much larger. This indicates that the link-capacity based inequalities (28), (29), (32), (33), (35), and (36) can indeed significantly strengthening the FP problem only with the connectivity based inequalities (18) and (20). In addition, with the increasing of the number of services, the gap improvement of (FP-II) becomes larger. This can be explained as follows. As the number of services increases, the left-hand sides of inequalities (32) and (33) become larger, thereby making inequalities (32) and (33) tighter. Consequently, the gap improvement achieved by the link-capacity based inequalities is also likely to be larger.
@@
Next, we illustrate the efficiency of the proposed connectivity based inequalities (18) and (20) and link-capacity based inequalities (28), (29), (32), (33), (35), and (36) for accelerating the convergence of the CBD algorithm (called CBD). Figs. 6 and 7 plot the average numbers of iterations and CPU times needed by the proposed CBDs with relaxations (FP), (FP-I), and (FP-II), respectively, to terminate. As observed, both of the derived families of valid inequalities can improve the performance of the proposed CBD. In particular, with the two proposed inequalities, the average number of iterations needed for the convergence of CBD is much smaller (less than $2$); and thus the CPU time taken by CBD is also much smaller.
@@
From the above computational results, we can conclude that the proposed CBD with relaxation (FP-II) (i.e., the FP problem with the two families of valid inequalities) significantly outperforms that with relaxation (FP) (i.e., the FP problem without the two families of valid inequalities). Due to this, we shall only test and compare the proposed CBD with relaxation (FP-II) with the other existing algorithms in the following.
V-B Comparison of CBD with SOTA Algorithms
@@
Figure 8: The number of feasible problem instances solved by CBDs with $\text{IterMax}=5,\,+\infty$, LPoR, LPdR, PSUM-R, and MBLP-S.
@@
Figure 9: The average objective value returned by CBDs with $\text{IterMax}=5,\,+\infty$, LPoR, LPdR, PSUM-R, and MBLP-S.
@@
Figure 10: The average CPU time returned by CBDs with $\text{IterMax}=5,\,+\infty$, LPoR, LPdR, PSUM-R, and MBLP-S.
@@
In this subsection, we compare the performance of the proposed CBD algorithm with SOTA algorithms such as the exact approach using standard MBLP solvers (called MBLP-S) in [7], the LP one-shot rounding (LPoR) algorithm in [16], the LP dynamic rounding (LPdR) algorithm in [17], and the penalty successive upper bound minimization with rounding (PSUM-R) algorithm in [18].
@@
Figs. 8 and 9 plot the number of feasible problem instances and the average objective value returned by the proposed CBDs with $\text{IterMax}=5,\,+\infty$, LPoR, LPdR, PSUM-R, and MBLP-S, respectively. When setting $\text{IterMax}=5$, the proposed CBD cannot be guaranteed to find an optimal solution of problem (NS), and the red-diamond curve in Fig. 8 is obtained as follows: for a problem instance, if the proposed CBD can find an optimal solution of problem (NS) within 5 iterations, it is counted as a feasible instance; otherwise, it is infeasible. From Figs. 8 and 9, we can observe the effectiveness of the proposed CBDs over LPoR, LPdR, and PSUM-R. In particular, compared with LPoR, LPdR, and PSUM-R, the proposed CBDs can find a feasible solution for much more problem instances (especially for instances with a large number of services) and return a much better solution with a much smaller objective value. The latter is reasonable, as stated in Section III-B, CBDs must return an optimal solution as long as they find a feasible solution. Another simulation finding is that with the increasing of IterMax, CBD can return a feasible solution for more problem instances. Nevertheless, when $\text{IterMax}=5$, CBD can return a feasible solution for most instances (as CBD with $\text{IterMax}=+\infty$ is able to find all truly feasible problem instances and the difference between the numbers of feasible problem instances solved by CBDs with $\text{IterMax}=5$ and $\text{IterMax}=+\infty$ is small in Fig. 8). This shows that in most cases, CBD can return an optimal solution within a few number of iterations, thanks to the two derived families of valid inequalities.
@@
The comparison of the solution efficiency of the proposed CBDs with $\text{IterMax}=5,\,+\infty$, LPoR, LPdR, PSUM-R, and MBLP-S is plotted in Fig. 10. We can observe from Fig. 10 that though both MBLP-S and the proposed CBD with $\text{IterMax}=+\infty$ can return an optimal solution (if the problem is feasible), CBD with $\text{IterMax}=+\infty$ is much more computationally efficient than MBLP-S. In particular, the average CPU times taken by CBD with $\text{IterMax}=+\infty$ are less than $0.7$ seconds, while the CPU times taken by MBLP-S in the large-scale cases (i.e., $|\mathcal{K}|\geq 13$) are generally larger than $5$ seconds. The solution efficiency of the proposed CBD with $\text{IterMax}=5$, LPoR, LPdR, and PSUM-R is comparable. In general, LPoR performs the best in terms of the CPU time, followed by CBD with $\text{IterMax}=5$ , LPdR, and PSUM-R.
@@
In summary, our simulation results show both of the effectiveness and efficiency of the proposed CBD. More specifically, compared with the LPoR, LPdR, and PSUM-R algorithms in [16], [17] and [18], the proposed CBD is able to find a much better solution (due to the information feedback between the two subproblems); compared with the exact approach in [7], the proposed CBD is significantly more computationally efficient, which is because of its decomposition nature.
@@ VI Conclusions
In this paper, we have proposed an efficient CBD algorithm for solving large-scale NS problems. The proposed algorithm decomposes the large-scale NS problem into two relatively easy FP and TR subproblems and solves them in an iterative fashion enabling a useful information feedback between them (to deal with the coupled constraints) until an optimal solution of the problem is found. To further improve the efficiency of the proposed algorithm, we have also developed two families of valid inequalities to speed up its convergence by judiciously taking the special structure of the considered problem into account. Two key features of the proposed algorithm, which make it particularly suitable to solve the large-scale NS problems, are as follows. First, it is guaranteed to find a globally optimal solution of the NS problem if the number of iterations between the two subproblems is allowed to be sufficiently large and it can return a much better solution than existing two-stage heuristic algorithms even though the number of iterations between the subproblems is small (e.g., 5). Second, the FP and TR subproblems in the proposed algorithm are a small MBLP problem and a polynomial time solvable LP problem, respectively, which are much easier to solve than the original problem. Simulation results show the effectiveness and efficiency of the proposed algorithm over the existing SODA algorithms in [7, 16, 17, 18]. In particular, when compared with algorithms based on LP relaxations [16, 17, 18], the proposed algorithm is able to find a much better solution; when compared with the algorithm needed to call MBLP solvers [7], the proposed algorithm is much more computationally efficient.
As the future work, it would be interesting to extend the proposed CBD algorithm to solve the virtual network embedding (VNE) problem in [2, 38], which has wide applications in wireless and optical communication networks. The VNE problem in [2, 38] is a variant of the NS problem that requires the traffic flow between the two nodes hosting two adjacent functions to be routed via a single path. To develop a CBD algorithm to solve the VNE problem, we need to judiciously design an algorithm for solving the traffic routing subproblem (which is an integer programming problem), enabling to find a valid inequality to cut off the current infeasible solution.
@@ Appendix A Proof of Theorem 3
$x_{v_{0}}^{k,s_{0}}+x_{v}^{k,s_{0}+1}\leq 1,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{~{}\forall~{}v\in\mathcal{V}(v_{0}),~{}\forall~{}v_{0}\in\mathcal{V}\backslash\mathcal{V}(S^{k})}}.$ (45)
@@
Next, we show that (18) holds for all $k\in\mathcal{K}$, $v\in\mathcal{V}(D^{k})$, and $s\in\mathcal{F}^{k}$ at $\bm{x}\in\mathcal{X}^{1}$ by an induction of $s$. By (16), $x_{v}^{k,s}=0$ holds for $s=\ell_{k}$. Suppose that $x_{v}^{k,s}=0$ holds for $s=s_{0},s_{0}+1,s_{0}+2,~{}\ldots,\ell_{k}$, where $1<s_{0}\leq\ell_{k}$. Then we have $x_{v}^{k,s_{0}}=0$ for all $v\in\mathcal{V}(D^{k})$. This, together with (1), implies
@@
$x_{v_{0}}^{k,s_{0}-1}+x_{v}^{k,s_{0}}\leq 1,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{~{}\forall~{}v\in\mathcal{V}(v_{0}),~{}\forall~{}v_{0}\in\mathcal{V}(D^{k})}}.$ (48)